OpenStudy (anonymous):

4x-3/7<3/8-x I started someone help me finish it

OpenStudy (anonymous):

\[\frac{ 4x-3 }{ 7 }<7(\frac{ 3 }{ 8 }-x)\]

OpenStudy (anonymous):

i multiplied each side by 7

OpenStudy (anonymous):

and got 4x-3<21/8-7x

OpenStudy (anonymous):

now i dont know what to do after this

OpenStudy (anonymous):

\[4x-3<\frac{ 21 }{ 8 }-7x\]

OpenStudy (ranga):

Isolate x. Add 7x to both sides.

OpenStudy (anonymous):

why not 4

OpenStudy (ranga):

You can do that too. With 4x you have to subtract it from both sides. With 7x you can dd to both sides.

OpenStudy (anonymous):

i would get the same answer right. Which ever side i used. And it can be with any problem???

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

thanks you just cleared everything up for me. I get confuse a lot

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

now do add 3 on both sides

OpenStudy (anonymous):

i have 11x-3<21/8 do i add 3

OpenStudy (ranga):

Yes add 3 to both sides

OpenStudy (anonymous):

i got 11<5

OpenStudy (anonymous):

wait sorry not that

OpenStudy (anonymous):

\[11x<5\frac{ 5 }{ 8 }\]

OpenStudy (ranga):

make the right side a full fraction than a mixed number.

OpenStudy (anonymous):

how do i do that

OpenStudy (ranga):

5 + 5/8 = (40 + 5)/8 = 45/8

OpenStudy (anonymous):

i got 45/88

OpenStudy (anonymous):

is this correct

OpenStudy (ranga):

Yes. But it should be x < 45/88